# ROScopter: A Multirotor Autopilot based on ROSflight 2.0

### 05 Mar 2026 v1

@@ II Related Work
Both PX4 and ArduPilot support many different airframes and vehicle configurations. However, when non-standard aircraft (e.g. tilt-rotor) are desired, many researchers opt to write firmware from scratch [4]. To avoid writing custom firmware from scratch for novel controllers, the authors in [4] develop a template to more easily integrate custom controllers into PX4. This template approach can reduce researcher effort when integrating custom controllers, but does not reduce the complexity or learning curve of the autopilot. ROScopter offers a reduced feature set, that reduces the size of the code-base and improves understandability of the autopilot.
@@
Other autopilots like [14] or [6] have large hobbyist communities, but are less suited to UAV research due limited support for onboard companion computers. Additionally, these projects do not have any official support for ROS 2.
@@ III-B1 State Propagation
The state estimator is a continuous-discrete formulation as in [3], meaning that the derivation and analysis utilizes the continuous-time equations that are then developed into discrete-time equations for use in the estimator. This has the advantage of having a direct connection to the familiar dynamical equations and a clear relationship to the discrete time formulation that follows.
The dynamics of the aircraft state are modeled as a function of the state and the input to the system $\boldsymbol{u}=\begin{bmatrix}\boldsymbol{a}^{\top}&\boldsymbol{\omega}^{\top}\end{bmatrix}^{\top}$.
@@
Since the actual inputs $\boldsymbol{a}$ and $\boldsymbol{\omega}$ are unavailable, we use the IMU to approximate these as $\boldsymbol{a}\approx\boldsymbol{y}_{\text{accel}}$ and $\boldsymbol{\omega}\approx\boldsymbol{y}_{\text{gyro}}-\boldsymbol{b}_{\text{gyro}}$. These are assumed to be the instantaneous measurement of the acceleration and angular rates and are held constant over the time step between estimation. The rotation matrix $R_{b}^{i}(\boldsymbol{\theta})$ expresses vectors in the body-frame in the inertial frame according to the current attitude estimate $\boldsymbol{\theta}$.
@@
is the relation from body rates to Euler angle rates. It properly accounts for the sequential frames that Euler angles are described in.
Note that $A$ is the continuous time Jacobian not the discrete time Jacobian. To find the discrete time Jacobian so it can be used in the EKF’s discrete updates, we use a second order approximation of the matrix exponential, $e^{AT_{s}}$, where $T_{s}$ is the period between state propagation updates in the EKF. It is given by
@@ III-B2 Measurement Updates
The measurements obtained by the sensors, $\boldsymbol{z}$, are incorporated into the estimate of the state $\boldsymbol{x}$ using the Kalman update based on the predicted measurement $h(\boldsymbol{x},\boldsymbol{u})$,
$\boldsymbol{x}_{k+1}=\boldsymbol{x}_{k}+K(z-h(\boldsymbol{x},\boldsymbol{u}))$ (8)
@@
If sensor measurement $z$ is a function of the states and the sensor reading the covariance of the innovation, $S$, will have more terms, in which case,
@@
Each measurement model, $h$, and observation Jacobian, $C$, for each sensor are given below.
@@
This finds the components of the magnetic field in the horizontal plane by projection, then finds the angle of that projection relative to true north. The magnetic field measurements need to be expressed in the inertial frame yawed by the heading of the aircraft. This allows for the calculation of the angle of the north axis using atan2, while compensating for the roll and pitch tilt of the aircraft.
This yields the simple measurement model,
@@
The contribution of the uncertainty of the states used to calculate $z_{\text{mag}}$, is found by calculating $F=\frac{\partial z}{\partial y}$ and $G=\frac{\partial z}{\partial\boldsymbol{x}}$ and using equation 12.
@@
The GNSS altitude measurement is omitted because commercially available single-band GNSS antennas can have large drifts in altitude measurement. The raw latitude and longitude angles of the GNSS are converted to a local north and east position relative to an initial latitude and longitude recorded on initialization, or provided through configuration. This is done by making a spherical earth assumption,
@@
where $r_{\text{earth}}$ is the radius of the Earth. These make up the measurement $y_{\text{gnss}}$.
The measurement model used for GNSS is,
$h_{\text{gnss}}=\begin{bmatrix}p_{n}\\ p_{e}\\ R_{b}^{i}(\boldsymbol{\theta})\boldsymbol{v}\end{bmatrix}$. (22)
This takes the estimated north and east positions directly from the state and the body velocities expressed in the inertial frame according to the current attitude estimate. This allows for tight coupling between the attitude angles and the high certainty velocity measurements from GNSS. This also tightly couples the gyroscope biases and the GNSS velocities.
This measurement model yields the observation Jacobian,
@@ III-C Navigation Stack
ROScopter’s navigation stack consists of a path planner, a path manager, and a path follower, as shown in Figure 1. The path planner is responsible for collecting waypoints and publishing them to the path manager. In the default implementation of ROScopter, the $i$th waypoint $w_{i}=(p_{i},\psi_{i})$ is defined by the user and is specified by a desired 3-D position $p_{i}$ and heading $\psi_{i}$.
@@
$u_{\text{traj}}(\tau)=[p(\tau),\dot{p}(\tau),\ddot{p}(\tau),\psi(\tau),\dot{\psi}(\tau),\ddot{\psi}(\tau)]^{T},$ (24)
where $u_{\text{traj}}$ is the desired trajectory. The path manager publishes each desired trajectory setpoint, $u_{\text{traj}}(\tau)$, at a configurable rate.
Here, $\tau\in[0,1]$ is normalized time, given by
@@
where $t$ is the time along the current waypoint leg, and $T_{\text{path}}$ is the total time it will take to travel the waypoint leg.
In the default implementation of ROScopter, the path manager generates straight-line trajectories between consecutive waypoints $p_{i}$ and $p_{i+1}$. Thus, $p(\tau)$ and $\psi(\tau)$ are linear interpolations between the desired positions and headings at consecutive waypoints, where $p(0)=p_{i}$, $\psi(0)=\psi_{i}$, $p(1)=p_{i+1}$, and $\psi(1)=\psi_{i+1}$. The rate of interpolation is given by the path parameter $\sigma(\tau)$, which can be any monotonically increasing function with $\sigma(0)=0$ and $\sigma(1)=1$. In ROScopter, we define $\sigma(\tau)$ to be a 5th-order smoothstep function. This results in zero velocity and acceleration at each waypoint, resulting in a smooth transition between waypoint legs.
The trajectory follower is a simple PID controller on position that incorporates position, velocity, acceleration feedforward terms from the differentially flat formulation of [5] to follow the trajectory setpoints from Eq. 24. As in [5], the trajectory follower module produces control setpoints of the form
$u_{\text{angle}}=[\phi^{d},\theta^{d},r^{d},T^{d}]^{T},$ (26)
where $u_{\text{angle}}$ is the desired setpoint published to an angle controller and is composed of the desired roll $\phi^{d}$, pitch $\theta^{d}$, yaw rate $r^{d}$, and thrust $T^{d}$, respectively. These reference setpoints are then published to the controller.
@@ III-D Controller
The ROScopter controller is responsible for taking in high-level controller commands and converting them to lower-level commands. Because different applications that feed into the controller produce different reference commands, the ROScopter controller architecture cascades as shown in Figure 2. Each entry point is a separate PID controller and is is described in Table I. In this formulation, users can send reference commands to any of the entry points in the ROScopter controller, and the commands cascade down to the lowest-level control loops. For example, under a particular navigation scheme, a user may send reference inertial velocities to the ROScopter controller. This velocity loop is closed by controller 3, which returns accelerations in the vehicle-1 frame (see [3]). These reference accelerations are passed to controller 2, and then to controller 6, where the resulting reference angle commands are published to the ROSflight firmware.
@@ 11 RollR PitchR YawR Thrust Roll rate, pitch rate, yaw rate, and thrust to pass-through
Many different types of control reference inputs are standard in multirotor applications, so this formulation of ROScopter prevents users from needing to write a low-level controller for all different types of standard reference commands. Additionally, a given input type is selected by setting the appropriate mode as each command message is published, so different control input types may be used arbitrarily. For example, the default configuration of ROScopter uses controller 4 during takeoff and controller 10 during other portions of flight. Controller 10 takes in reference setpoints computed by the trajectory follower, as shown in Eq 26.
ROScopter is designed to work with an FCU running ROSflight firmware and thus returns commands in one of the three modes that ROSflight accepts, as shown in Figure 2. As described in [10], the ROSflight firmware contains two low-level controllers, an angle controller and a rate controller. Additionally, ROSflight allows commands to bypass all low-level loops on the firmware and progress directly to the mixer, called pass-through mode. Thus, the ROScopter controller publishes one of the following command tuples
$u_{\text{angle}}$ $=[0,0,\delta_{t}^{d},\phi^{d},\theta^{d},r^{d},0_{1\times 4}]^{T},$ (27)
$u_{\text{rate}}$ $=[0,0,\delta_{t}^{d},p^{d},q^{d},r^{d},0_{1\times 4}]^{T},$ (28)
$u_{\text{pass-through}}$ $=[0,0,T_{z}^{d},Q_{x}^{d},Q_{y}^{d},Q_{z}^{d},0_{1\times 4}]^{T},$ (29)
where $\phi^{d}$, $\theta^{d}$, are the desired roll and pitch, $p^{d}$, $q^{d}$, $r^{d}$ are the desired roll rate, pitch rate, and yaw rate, $\delta_{t}^{d}\in[0,1]$ is the desired throttle setpoint, and $Q_{x}^{d}$, $Q_{y}^{d}$, $Q_{z}^{d}$, $T_{z}^{d}$ are the desired roll torque, pitch torque, yaw torque, and thrust in the body $z$ direction, respectively. The zeros in these command messages is to comply with the default interpretation of this $u$ command vector, described in [10].
@@ IV-A Default ROScopter behavior
Although this simple functionality will not satisfy all researchers’ needs, the default ROScopter autonomy stack is useful in many applications. Applications such as optimal search-and-rescue path planning may only require that a multirotor travel to a particular sequence of waypoints; in this case, ROScopter could effectively sit underneath application code.
@@ IV-B Customizing ROScopter
Every module in Figure 1 is implemented as a ROS 2 node with clearly defined ROS 2 interfaces (e.g. publishers, subscribers, and service calls). Additionally, as discussed in Section III-A, each module is implemented using an inheritance pattern to facilitate modification. This reduces dependencies between modules, isolates functionality, and allows for easy customization of a particular module without modifying any other modules, thus lowering the barrier to entry for researchers. For example, a different state estimator with different measurement or update functions could be implemented by rewriting the existing estimator module or by sub-classing the estimator interface. As long as the ROS 2 interfaces between the estimator and all other modules remain the same (in this case, the state estimate message), then the other modules are agnostic to how the state estimate values are computed, and the new estimator code will integrate seamlessly into ROScopter.
@@ IV-C Intended workflow for ROScopter
When using the ROScopter autnonomy stack, users should first determine which nodes should be replaced or modified given a particular application. Application code can then be developed and integrated into the autonomy stack as described in the previous sections. Simulation tests can easily be run by configuring the autonomy stack at runtime and running application code instead of the default ROScopter modules.
@@ V-B Comparison to PX4
To compare the closed-loop controller response for waypoint-following missions, a Pixracer Pro FCU was flashed with ROSflight firmware and mounted to a HolyBro x650 frame. For these experiments, a Raspberry Pi 5 was used as the companion computer (configuration 1 as described in [10]. A mission was loaded to the ROScopter path planner and was flown. Then, the Pixracer Pro was flashed with PX4 firmware version v1.15.4 and mounted to the same HolyBro x650 frame. The same mission was loaded using a laptop with the QGroundControl ground control software, and was flown. Figure 4 shows ROScopter’s closed-loop performance when following the same trajectory shown in Figure 3. Figure 6 shows PX4’s closed-loop performance when following this same trajectory. Table III shows the root-mean-squared errors between each autopilot’s estimated state and the ideal straight-line path between waypoints. We expect that both responses could be improved with more rigorous gain tuning. While PX4 contains many additional features, excellent autopilot functionality, and improved performance, these results show that ROScopter offers similar performance for basic waypoint-following functionality while maintaining a lean, understandable, modular, and clean code-base.

### 06 Mar 2026 v2

@@ II Related Work
Both PX4 and ArduPilot support many different airframes and vehicle configurations. However, when non-standard aircraft (e.g. tilt-rotor) are desired, many researchers opt to write firmware from scratch [4]. To avoid writing custom firmware from scratch for novel controllers, the authors in [4] develop a template to more easily integrate custom controllers into PX4. This template approach can reduce researcher effort when integrating custom controllers, but does not reduce the complexity or learning curve of the autopilot. ROScopter offers a reduced feature set, that decreases the size of the code-base and improves understandability of the autopilot.
@@
Other autopilots like [14] or [6] have large hobbyist communities, but are less suited to UAV research due to limited support for onboard companion computers. Additionally, these projects do not have any official support for ROS 2.
@@ III-B1 State Propagation
The state estimator is a continuous-discrete formulation as in [3]. This has the advantage of having a direct connection to the familiar dynamical equations and a clear relationship to the discrete time formulation that follows.
The dynamics of the aircraft state are modeled as a function of the state and the input to the system $\boldsymbol{u}=\begin{bmatrix}\boldsymbol{a}^{\top}&\boldsymbol{\omega}^{\top}\end{bmatrix}^{\top}$, where $\boldsymbol{a}$ is specific acceleration and $\boldsymbol{\omega}$ is the angular velocity.
@@
Since the actual inputs $\boldsymbol{a}$ and $\boldsymbol{\omega}$ are unavailable, we use the IMU to approximate these as $\boldsymbol{a}\approx\boldsymbol{y}_{\text{accel}}$ and $\boldsymbol{\omega}\approx\boldsymbol{y}_{\text{gyro}}-\boldsymbol{b}_{\text{gyro}}$. These are assumed to be the instantaneous measurement of the specific acceleration and angular rates and are held constant over the time step between estimates. The rotation matrix $R_{b}^{i}(\boldsymbol{\theta})$ transforms vectors in the body-frame into the inertial frame according to the current attitude estimate $\boldsymbol{\theta}$.
@@
is the relation from body rates to Euler angle rates and properly accounts for the sequential frames associated with the Euler angles.
Note that $A$ is the continuous-time Jacobian and not the discrete-time Jacobian. To find the discrete-time Jacobian so it can be used in the EKF’s discrete updates, we use a second order approximation of the matrix exponential $e^{AT_{s}}$, where $T_{s}$ is the period between state propagation updates in the EKF and is given by
@@ III-B2 Measurement Updates
The measurements $\boldsymbol{z}$, are incorporated into the estimate of the state $\boldsymbol{x}$ using the Kalman update based on the predicted measurement $h(\boldsymbol{x},\boldsymbol{u})$:
$\boldsymbol{x}_{k+1}=\boldsymbol{x}_{k}+K(z-h(\boldsymbol{x}_{k},\boldsymbol{u}_{k}))$ (8)
@@
If sensor measurement $z$ is a function of the states and the sensor reading then the covariance of the innovation $S$ will have more terms, in which case,
@@
The measurement model $h$, and observation Jacobian $C$, for each sensor are given below.
@@
Eq. (15) finds the components of the magnetic field $\boldsymbol{m}$ in the horizontal plane by projection using $R_{b}^{v_{1}}$ to express the magnetic field measured in the yawed inertial frame $v_{1}$, and then finds the angle of that projection relative to true north.
The measurement model is simply
@@
The contribution of the uncertainty of the states used to calculate $z_{\text{mag}}$, is found by calculating $F=\frac{\partial z}{\partial y}$ and $G=\frac{\partial z}{\partial\boldsymbol{x}}$ and using Eq. (12).
@@
The GNSS altitude measurement is omitted because commercially available single-band GNSS antennas can have large drifts in altitude measurement. The raw latitude and longitude angles of the GNSS are converted to a local north and east position relative to an initial latitude and longitude recorded on initialization, or provided through configuration. This is done by making a spherical earth assumption:
@@
where $r_{\text{earth}}$ is the radius of the Earth.
The measurement model used for GNSS is
$h_{\text{gnss}}=\begin{bmatrix}p_{n}\\ p_{e}\\ R_{b}^{i}(\boldsymbol{\theta})\boldsymbol{v}\end{bmatrix}$, (22)
taking the estimated north and east positions directly from the state and the body velocities expressed in the inertial frame according to the current attitude estimate. This allows for tight coupling between the attitude angles and the high certainty velocity measurements from GNSS. An additional advantage is that the gyroscope biases and the GNSS velocities are tightly coupled.
The measurement model yields the observation Jacobian
@@ III-C Navigation Stack
ROScopter’s navigation stack consists of a path planner, a path manager, and a path follower, as shown in Figure 1. The path planner is responsible for collecting waypoints and publishing them to the path manager. In the default implementation of ROScopter, the $i$th waypoint $\boldsymbol{w}_{i}=(\boldsymbol{p}_{i},\psi_{i})$ is defined by the user and is specified by a desired 3-D position $\boldsymbol{p}_{i}$ and heading $\psi_{i}$.
@@
$\boldsymbol{u}_{\text{traj}}(\tau)=[\boldsymbol{p}(\tau),\dot{\boldsymbol{p}}(\tau),\ddot{\boldsymbol{p}}(\tau),\psi(\tau),\dot{\psi}(\tau),\ddot{\psi}(\tau)]^{T},$ (24)
where $\boldsymbol{u}_{\text{traj}}$ is the desired trajectory. The path manager publishes each desired trajectory setpoint, $\boldsymbol{u}_{\text{traj}}(\tau)$, at a configurable rate, where $\tau\in[0,1]$ is normalized time, given by
@@
and where $t$ is the time along the current waypoint leg, and $T_{\text{path}}$ is the total time it will take to travel the waypoint leg.
In the default implementation of ROScopter, the path manager generates straight-line trajectories between consecutive waypoints $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{i+1}$. Thus, $\boldsymbol{p}(\tau)$ and $\psi(\tau)$ are linear interpolations between the desired positions and headings at consecutive waypoints, where $\boldsymbol{p}(0)=\boldsymbol{p}_{i}$, $\psi(0)=\psi_{i}$, $\boldsymbol{p}(1)=\boldsymbol{p}_{i+1}$, and $\psi(1)=\psi_{i+1}$. The rate of interpolation is given by the path parameter $\sigma(\tau)$, which can be any monotonically increasing function with $\sigma(0)=0$ and $\sigma(1)=1$. In ROScopter, we define $\sigma(\tau)$ to be a 5th-order smoothstep function, resulting in zero velocity and acceleration at each waypoint, giving a smooth transition between waypoint legs.
The trajectory follower is a simple PID controller on position that incorporates position, velocity, acceleration feedforward terms from the differentially flat formulation of [5] to follow the trajectory setpoints from Eq. (24). As in [5], the trajectory follower module produces control setpoints of the form
$\boldsymbol{u}_{\text{angle}}=[\phi^{d},\theta^{d},r^{d},T^{d}]^{T},$ (26)
where $\boldsymbol{u}_{\text{angle}}$ is the desired setpoint published to an angle controller and is composed of the desired roll $\phi^{d}$, pitch $\theta^{d}$, yaw rate $r^{d}$, and thrust $T^{d}$, respectively. These reference setpoints are then published to the controller.
@@ III-D Controller
The ROScopter controller is responsible for taking in upstream controller commands and converting them to low-level commands. Because different applications that feed into the controller (e.g. the ROScopter trajectory follower) produce different reference commands, the ROScopter controller allows users to select a variety of input reference setpoint types as shown in Figure 2. Each box shown in Figure 2 is a separate PID controller and is described in Table I. In this formulation, users can send reference commands to any of the entry points in the ROScopter controller, and the commands cascade down to the lowest-level control loops. For example, under a particular navigation scheme, a user may send reference inertial velocities to the ROScopter controller. This velocity loop is closed by controller 3, which returns accelerations in the vehicle-1 frame (see [3]). These reference accelerations are passed to controller 2, and then to controller 6, where the resulting reference angle commands are published to the ROSflight firmware.
@@ 11 RollR PitchR YawR Thrust Roll rate, pitch rate, yaw rate, and thrust to pass-through
Many different types of control reference inputs are standard in multirotor applications, so this formulation of ROScopter prevents users from needing to write a low-level controller for all different types of standard reference commands. Additionally, a given input type is selected by setting the appropriate mode as each command message is published, so different control input types may be used arbitrarily. For example, the default configuration of ROScopter uses controller 4 during takeoff and controller 10 during other portions of flight. Controller 10 takes in reference setpoints computed by the trajectory follower, as shown in Eq. (26).
ROScopter is designed to work with a flight controller unit (FCU) running ROSflight firmware and thus returns commands in one of the three modes that ROSflight accepts, as shown in Figure 2. As described in [10], the ROSflight firmware contains two low-level controllers, an angle controller and a rate controller. Additionally, ROSflight allows commands to bypass all low-level loops on the firmware and progress directly to the mixer, called pass-through mode. Thus, the ROScopter controller publishes one of the following command vectors
$\boldsymbol{u}_{\text{angle}}$ $=[0,0,\delta_{t}^{d},\phi^{d},\theta^{d},r^{d},\boldsymbol{0}_{1\times 4}]^{T},$ (27)
$\boldsymbol{u}_{\text{rate}}$ $=[0,0,\delta_{t}^{d},p^{d},q^{d},r^{d},\boldsymbol{0}_{1\times 4}]^{T},$ (28)
$\boldsymbol{u}_{\text{pass-through}}$ $=[0,0,T_{z}^{d},Q_{x}^{d},Q_{y}^{d},Q_{z}^{d},\boldsymbol{0}_{1\times 4}]^{T},$ (29)
where $\phi^{d}$, $\theta^{d}$, are the desired roll and pitch, $p^{d}$, $q^{d}$, $r^{d}$ are the desired roll rate, pitch rate, and yaw rate, $\delta_{t}^{d}\in[0,1]$ is the desired throttle setpoint, and $Q_{x}^{d}$, $Q_{y}^{d}$, $Q_{z}^{d}$, $T_{z}^{d}$ are the desired roll torque, pitch torque, yaw torque, and thrust in the body $z$ direction, respectively. The zeros in these command messages are to comply with the default interpretation of this $\boldsymbol{u}$ command vector, described in [10].
@@ IV-A Default ROScopter behavior
Although this simple functionality will not satisfy all researchers’ needs, the default ROScopter autonomy stack is useful in many applications. Applications such as search-and-rescue may only require that a multirotor travel to a particular sequence of waypoints; in this case, ROScopter could effectively sit underneath application code.
@@ IV-B Customizing ROScopter
Every module in Figure 1 is implemented as a ROS 2 node with clearly defined ROS 2 interfaces (e.g. publishers, subscribers, and service calls). Additionally, as discussed in Section III-A, each module is implemented using an inheritance pattern to facilitate modification. This reduces dependencies between modules, isolates functionality, and allows for easy customization of a particular module without modifying any other modules, thus lowering the barrier to entry. For example, a different state estimator with different measurement or update functions could be implemented by rewriting the existing estimator module or by sub-classing the estimator interface. As long as the ROS 2 interfaces between the estimator and all other modules remain the same (in this case, the state estimate message), then the other modules are agnostic to how the state estimate values are computed, and the new estimator code will integrate seamlessly into ROScopter.
@@ IV-C Intended workflow for ROScopter
When using the ROScopter autonomy stack, users should first determine which nodes should be replaced or modified given a particular application. Application code can then be developed and integrated into the autonomy stack as described in the previous sections. Simulation tests can easily be run by configuring the autonomy stack at runtime and running application code instead of the default ROScopter modules.
@@ V-B Comparison to PX4
To compare the closed-loop controller response for waypoint-following missions, a Pixracer Pro FCU was flashed with ROSflight firmware and mounted to a HolyBro x650 frame. For these experiments, a Raspberry Pi 5 was used as the companion computer (configuration 1 as described in [10]. A mission was loaded to the ROScopter path planner and was flown. The Pixracer Pro was flashed with PX4 firmware version v1.15.4 and mounted to the same HolyBro x650 frame. The same mission was loaded using a laptop with the QGroundControl software, and was flown. Figure 4 shows ROScopter’s closed-loop performance when following the same trajectory shown in Figure 3. Figure 6 shows PX4’s closed-loop performance when following this same trajectory. Table III shows the root-mean-squared errors between each autopilot’s estimated state and the ideal straight-line path between waypoints. We expect that both responses could be improved with more rigorous gain tuning. While PX4 contains many additional features, excellent autopilot functionality, and improved performance, these results show that ROScopter offers similar performance for basic waypoint-following functionality while maintaining a lean, understandable, modular, and clean code-base.